# Symplectic mapping class relations from pencil pairs

### v2

@@ 1. Introduction
In [O21] Oba finds a symplectic mapping class relation on a Weinstein $4$-manifold by applying the classification of del Pezzo surfaces [F80] to find a pencil pair. Our main application applies classification results for polarized $K3$ varieties [M87, SD74] and Fano $3$-folds [I77, I77, MM82] to find new pencil pairs, and so new examples of mapping class relations on Weinstein $4$-manifolds. Here is one such family of examples.
@@ 1.1. Conventions
All algebraic varieties in this article are smooth and defined over $\mathbb{C}$. Throughout $\dim$ and $\dim_{\mathbb{C}}$ refer to real and complex dimension, respectively. On a closed manifold $X$ we use $\operatorname{PD}_{X}$ for Poincaré duality in either direction $H_{k}(X)\leftrightarrow H^{\dim X-k}(X)$. Cotangent disk bundles (determined by some Riemannian metric) over a smooth manifold $X$ will be denoted $\mathbb{D}^{\ast}X\subset T^{\ast}X$. On $T^{\ast}X$, the canonical $1$-form is denoted $pdq$ whose differential, denoted $dp\wedge dq$ is the canonical symplectic form. All other symplectic manifolds in this article will be complex projective varieties $X,Z,B,\mathcal{M}\subset\mathbb{CP}^{n}$ or affine varieties $F,W\subset\mathbb{C}^{n}$, with their symplectic structures inherited from the ambient space. For a function $h$ on a symplectic manifold with symplectic form $\omega$, the Hamiltonian vector field $X_{h}$ is defined by $dh=\omega(\ast,X_{h})$.
@@ 2.1. Symplectic mapping class relations
where $b=b(\left\lVert p\right\rVert)$ is a function which equals 00 near $\left\lVert p\right\rVert=0$ and $\pi$ near $\left\lVert p\right\rVert=1$, $X_{b}$ is the Hamiltonian vector field, and $(-\operatorname{Id}_{\mathbb{S}^{n}})^{\ast}$ is the symplectomorphism induced by the antipodal map $q\mapsto-q$ on $\mathbb{S}^{n}$. In words, $\operatorname{Flow}^{1}_{X_{b}}$ applies the time $b(\left\lVert p\right\rVert)$ geodesic flow to (co)vectors of length $\left\lVert p\right\rVert$. In particular, when $\left\lVert p\right\rVert=1$, the time $1$-flow will send a vector over $q\in\mathbb{S}^{n}$ to a vector over $-q$ and the $(-\operatorname{Id}_{\mathbb{S}^{n}})^{\ast}$ correction ensures that $\tau_{L}$ is the identity near $\partial\mathbb{D}^{\ast}\mathbb{S}^{n}$. Extend $\tau_{L}$ to $F\setminus N_{F}(L)$ as the identity. The dependence on $b$ is irrelevant when we view $\tau_{L}\in\pi_{0}\operatorname{Symp}^{c}$. It can be seen by drawing pictures that when $\dim F=2$, $\tau_{L}$ coincides with the usual notion of a (right-handed) Dehn twist along a simple closed curve on a Riemann surface. The action of $\tau_{L}$ on $H_{\ast}(F)$ is given by the Picard-Lefschetz formula
(2.1.1) $(\tau_{L})h=\begin{cases}h+(-1)^{\frac{1}{2}(n+1)(n+2)}(h\bullet[L])[L]&h\in H_{n}(F),\\ h&\text{otherwise.}\\ \end{cases}$
@@ 2.2. Known relations and counting twists
    $[L][L^{\prime}][L]^{-1}[L^{\prime}]^{-1}$ when the classes have representatives $L,L^{\prime}$ which are disjoint,
@@
    $[L][L^{\prime}][L]^{-1}[\tau_{L}L^{\prime}]^{-1}$ for any $L,L^{\prime}$,
@@
    and $[L][L^{\prime}][L][L^{\prime}]^{-1}[L]^{-1}[L^{\prime}]^{-1}$ when the classes have representatives $L,L^{\prime}$ intersecting transversely in $F$ a single point.
@@
$[\vec{L}]=[L_{k}]\cdots[L_{1}]$
which can be viewed as an element of either $T$ or $T/B$. A positive relation $\vec{L},\vec{L}^{\prime}$ is non-braid-like if $[\vec{L}]\neq[\vec{L}^{\prime}]\in T/B$. It’s clear that if $|\vec{L}|\neq|\vec{L}^{\prime}|$, then the relation is non-braid-like. Examples of non-braid-like relations include the lantern and chain relations on punctured Riemann surfaces, cf. [OS04b, §15].
@@
Consider the action of a single Dehn twist $\tau_{L}$ on the determinant line $\ell=\bigwedge^{top}H_{n}(F)$. Since $n$ is even, $[L]\bullet[L]=-2$ which isn’t divisible by a squared integer, implying that $[L]\in H_{n}(F)$ is primitive. Hence we can extend $[L]$ to a basis of $H_{n}(F)$ with respect to which $\tau_{L}\in\operatorname{Aut}(H_{n})$ can be expressed as a matrix. The first row has $-1$ in the first entry by Equation (2.1.2), with the rest of the entries being zeros. In the $j$th row for $j>1$, Equation (2.1.1) determines the first entry, the $j$th entry is $1$, and all other entries are zero. Therefore $\tau_{L}$ acts on $\ell$ by $-1$. For a general $\vec{L}\in T$, $\tau\vec{L}$ acts on the determinant line by $(-1)^{p\vec{L}}$. Since we’re assuming that $\tau\vec{L}$ acts trivially on $H_{n}(F)$, it must act trivially on $\ell$ and $2$ must divide $p\vec{L}$. ∎
@@ 2.3. Some contact-topological motivations
The example below demonstrates that the only constraints on the existence of a $(N,\xi)$ determining a fixed Euler number is unconstrained, beyond a cobordism-theoretic restriction that it must be even for $n$ even.
By [HH19, Theorem 1.2.3], any oriented hypersurface in a contact manifold $(M,\xi)$ of $\dim M\geq 5$ can be $\mathcal{C}^{0}$-approximated by a convex hypersurface $S\subset M$ whose $W$ and $W^{\prime}$ are Weinstein. Moreover [Br23, Theorem 1.8] states that such an $S$ can be assumed to be the result of the above construction determined by a positive relation. Combining [HH19] and [Br23] we see that convex hypersurfaces exist in abundance, as do positive relations. Moreover, Equation (2.3.1) tells us that there are lots of positive relations $\vec{L},\vec{L}^{\prime}$ which are not braid-like. Here are some examples.
@@
For a fixed $n\in\mathbb{Z}_{\geq 2}$ and a partition $\vec{n}=(n_{1},\dots,n_{m})$, $n_{i}\leq n_{i+1}$, $\sum n_{i}=n$, write
$\mathbb{CP}^{\vec{n}}=\mathbb{CP}^{n_{1}}\times\cdots\times\mathbb{CP}^{n_{m}},\quad M^{\vec{n}}=\mathbb{S}^{1}\times\mathbb{CP}^{\vec{n}}$
The complex structure on $T\mathbb{CP}^{\vec{n}}$ defines an almost contact structure on the $\dim=2n+1$ manifold $M^{\vec{n}}$. By [BEM15] – which includes a definition of almost contact structures – there is a (genuine) contact structure $\xi^{\vec{n}}$ on $M^{\vec{n}}$ which is homotopic to $T\mathbb{CP}^{\vec{n}}\subset TM^{\vec{n}}$ through almost contact structures.
Let $(M,\xi)$ be the contact connected sum of the $(M^{\vec{n}},\xi^{\vec{n}})$, over all partitions $\vec{n}$, the number of which we’ll write as $N_{n}$. For $\vec{r}=(r_{\vec{n}})\in\mathbb{Z}^{N_{n}}$ consider the homology class $h_{\vec{r}}\in H_{2n}(M)$ given by the sum of the $r_{\vec{n}}[\{\theta\}\times\mathbb{CP}^{\vec{n}}]$, with each summand living in the image of the composition
$H_{2n}(M^{\vec{n}})\xrightarrow{\simeq}H_{2n}(M^{\vec{n}}\setminus\mathbb{D}^{2n+1})\rightarrow H_{2n}(M).$
@@
It’s not difficult to see that $h_{\vec{r}}$ has a connected, embedded, oriented representative, $S_{0}$. By [HH19] we can apply a $\mathcal{C}^{0}$-small perturbation to $S_{0}$ to obtain an orientable, $\operatorname{codim}=1$ submanifold $S\subset M$ which is convex with neighborhood $(N,\xi)$. The contact structure $\xi|_{S}$ gives $S$ the structure of a stably almost complex manifold whose cobordism class is $[S]=\sum r_{\vec{n}}\mathbb{CP}^{\vec{n}}\in\Omega^{U}_{2n}$. Since the $\mathbb{CP}^{\vec{n}}$ additively generate $\Omega^{U}_{2n}$ (cf. [MiS74]), we see that any almost complex cobordism class can be represented by a convex hypersurface.
@@
By [Br23, Theorem 1.8] there is some Weinstein $(F,\beta_{F})$ of $\dim=2n-2$ having a positive relation $\vec{L},\vec{L}^{\prime}$ determining $(N,\xi)$. By Equation (2.3.1), $\int_{S}e(\xi)=\int_{S}c_{n}(\xi)$ determines the difference in the lengths of the positive factorizations. Note that the image of $\int_{\ast}c_{n}:\Omega^{U}_{2n}\rightarrow\mathbb{Z}$ is $2\mathbb{Z}$ when $n$ is odd and $\mathbb{Z}$ when $n$ is even. Taking $m=n-1$, we see that for every $m\geq 2$ there are Weinstein $(F,\beta_{F})$ of $\dim=2m$ having positive relations with $|\vec{L}|-|\vec{L}^{\prime}|$ being any integer when $m$ is odd and any even integer when $m$ is even.
@@
As convex hypersurfaces $S$ are only $\mathcal{C}^{0}$-dense, rather than $\mathcal{C}^{1}$-dense in $\dim\geq 5$ contact manifolds (cf. [Ch24]), it is very difficult to explicitly describe such convex surface neighborhoods $(N\supset S,\xi)$, even for convex hypersurfaces in well-studied $(M,\xi)$. In searching for symplectic mapping class relations, we seek to have some interesting examples of convex hypersurfaces at least for which the $(F,\beta_{F})$ can be described somewhat explicitly (e.g. as affine or quasi-projective varieties). In a follow-up article, we will study the $(N,\xi)$ determined by some of the positive relations in this article using the techniques of [Av23a].
@@ 2.4. Review of algebraic Lefschetz pencils
Let $\operatorname{\mathcal{L}}$ be a very ample line bundle on a smooth, closed variety $X$. By definition, this means that a $\mathbb{C}$-basis $(\boldsymbol{\mathfrak{s}}_{i})_{i=0}^{N}$ of $H^{0}(X,\operatorname{\mathcal{L}})$, the space of holomorphic sections of $\operatorname{\mathcal{L}}\to X$, determines a holomorphic embedding
@@
gives an embedding into $\mathbb{C}^{N}$. So each $X\setminus\{\boldsymbol{\mathfrak{s}}_{i}=0\}$ is naturally an affine variety. Replacing $\operatorname{\mathcal{L}}$ with $\operatorname{\mathcal{L}}^{\otimes k}=\phi_{\operatorname{\mathcal{L}}}^{\ast}\mathcal{O}_{k}$ for $k\in\mathbb{Z}_{>0}$ gives more very ample line bundles on $X$. Since $\phi_{\operatorname{\mathcal{L}}}$ is holomorphic,
@@
A pair $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ of distinct holomorphic sections of $\operatorname{\mathcal{L}}$ yields
@@
We say that $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ is a Lefschetz pencil for $\operatorname{\mathcal{L}}$ if the following criteria are met:
@@
The symplectomorphism type of the $(Z_{1},\omega_{\operatorname{\mathcal{L}}},B)$ associated to a pencil for $\operatorname{\mathcal{L}}$ is independent of sections $\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1}$.
@@
Consider a generic $T\in[0,1]$ family $(\boldsymbol{\mathfrak{s}}_{0}^{T},\boldsymbol{\mathfrak{s}}_{1}^{T})$ of pairs of sections of $\operatorname{\mathcal{L}}$. By Bertini’s theorem the set of such pairs for which $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ fails to be a Lefschetz-pencil has complex codimension $\geq 1$, that is, real codimension $\geq 2$. Therefore we can assume that the path of pairs of sections determines a path of pencils. Applying a Darboux-Moser-Weinstein argument [MS99, §3.2] to the $T$-family establishes the claim. ∎
@@
The pencil $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ determines a compactly supported symplectomorphism of $(F,d\beta)$ in two ways. One way to express this symplectomorphism comes from studying the monodromy of $\pi$. Let $\eta_{i}$ be a collection of embedded, oriented, $[0,1]$-parameterized paths in $\mathbb{C}$ indexed by the $z_{i}\in\operatorname{Critv}(\pi)$, beginning at $0\in\mathbb{C}$, ending at $z_{i}$, and disjoint away from $0\in\mathbb{C}$. We require that the indices $i$ are ordered so that tangent vectors of the $\eta_{i}$ at 00 are ordered counterclockwise. Each $\eta_{i}$ determines a Lagrangian disk $D_{i}$ in $W$ for which $\pi D_{i}=\eta$ and for which $L_{i}=\partial D_{i}$ is an exact Lagrangian sphere in $F$. The $\eta_{i}$ are vanishing paths and the $D_{i}$ are their associated Lefschetz thimbles which provide framings of the $L_{i}$. Let $\gamma\subset\mathbb{C}\setminus\operatorname{Critv}(\pi)$ be a parameterized simple closed curve based at 00 and oriented counterclockwise, otherwise disjoint from the $\eta_{i}$, encircling all of the $\eta_{i}$, and encircling $\eta_{1}$ first. Then the monodromy morphism $\Phi_{\gamma}\in\operatorname{Symp}^{c}(F,\omega_{\operatorname{\mathcal{L}}})$ can be expressed as a product of Dehn twists

### v3

@@ 1. Introduction
In [O21] Oba finds a symplectic mapping class relation on a Weinstein $4$-manifold by applying the classification of del Pezzo surfaces [F80] to find a pencil pair. Our main application leverages classification results for polarized $K3$ varieties [M87, SD74] and Fano $3$-folds [I77, I77, MM82] to find new pencil pairs, and so new examples of mapping class relations on Weinstein $4$-manifolds. Here is one such family of examples.
@@ 1.1. Conventions
All algebraic varieties in this article are smooth and defined over $\mathbb{C}$. Throughout $\dim$ and $\dim_{\mathbb{C}}$ refer to real and complex dimension, respectively. On a closed manifold $X$ we use $\operatorname{PD}_{X}$ for Poincaré duality in either direction $H_{k}(X)\leftrightarrow H^{\dim X-k}(X)$. Cotangent disk bundles (determined by some Riemannian metric) over a smooth manifold $X$ will be denoted $\mathbb{D}^{\ast}X\subset T^{\ast}X$. On $T^{\ast}X$, the canonical $1$-form is denoted pdqpdq whose differential, denoted $dp\wedge dq$ is the canonical symplectic form. All other symplectic manifolds in this article will be complex projective varieties $X,Z,B,\mathcal{M}\subset\mathbb{CP}^{n}$ or affine varieties $F,W\subset\mathbb{C}^{n}$, with their symplectic structures inherited from the ambient space. For a function $h$ on a symplectic manifold with symplectic form $\omega$, the Hamiltonian vector field $X_{h}$ is defined by $dh=\omega(\ast,X_{h})$.
@@ 2.1. Symplectic mapping class relations
where $b=b(\left\lVert p\right\rVert)$ is a function which equals 0 near $\left\lVert p\right\rVert=0$ and $\pi$ near $\left\lVert p\right\rVert=1$, $X_{b}$ is the Hamiltonian vector field, and $(-\operatorname{Id}_{\mathbb{S}^{n}})^{\ast}$ is the symplectomorphism induced by the antipodal map $q\mapsto-q$ on $\mathbb{S}^{n}$. In words, $\operatorname{Flow}^{1}_{X_{b}}$ applies the time $b(\left\lVert p\right\rVert)$ geodesic flow to (co)vectors of length $\left\lVert p\right\rVert$. In particular, when $\left\lVert p\right\rVert=1$, the time $1$-flow will send a vector over $q\in\mathbb{S}^{n}$ to a vector over q-q and the $(-\operatorname{Id}_{\mathbb{S}^{n}})^{\ast}$ correction ensures that $\tau_{L}$ is the identity near $\partial\mathbb{D}^{\ast}\mathbb{S}^{n}$. Extend $\tau_{L}$ to $F\setminus N_{F}(L)$ as the identity. The dependence on $b$ is irrelevant when we view $\tau_{L}\in\pi_{0}\operatorname{Symp}^{c}$. It can be seen by drawing pictures that when $\dim F=2$, $\tau_{L}$ coincides with the usual notion of a (right-handed) Dehn twist along a simple closed curve on a Riemann surface. The action of $\tau_{L}$ on $H_{\ast}(F)$ is given by the Picard-Lefschetz formula
[…]
@@ 2.2. Known relations and counting twists
    $L][L^{\prime}][L]^{-1}[L^{\prime}]^{-1}$ when the classes have representatives $L,L^{\prime}$ which are disjoint,
@@
    $L][L^{\prime}][L]^{-1}[\tau_{L}L^{\prime}]^{-1}$ for any $L,L^{\prime}$,
@@
    and $L][L^{\prime}][L][L^{\prime}]^{-1}[L]^{-1}[L^{\prime}]^{-1}$ when the classes have representatives $L,L^{\prime}$ intersecting transversely in $F$ a single point.
@@
$\vec{L}]=[L_{k}]\cdots[L_{1}]$
which can be viewed as an element of either $T$ or $T/B$. A positive relation $\vec{L},\vec{L}^{\prime}$ is non-braid-like if $\vec{L}]\neq[\vec{L}^{\prime}]\in T/B$. It’s clear that if $|\vec{L}|\neq|\vec{L}^{\prime}|$, then the relation is non-braid-like. Examples of non-braid-like relations include the lantern and chain relations on punctured Riemann surfaces, cf. [OS04b, §15].
@@
Consider the action of a single Dehn twist $\tau_{L}$ on the determinant line $\ell=\bigwedge^{top}H_{n}(F)$. Since $n$ is even, $L]\bullet[L]=-2$ which isn’t divisible by a squared integer, implying that $L]\in H_{n}(F)$ is primitive. Hence we can extend $[L]$ to a basis of $H_{n}(F)$ with respect to which $\tau_{L}\in\operatorname{Aut}(H_{n})$ can be expressed as a matrix. The first row has 1-1 in the first entry by Equation (2.1.2), with the rest of the entries being zeros. In the $j$th row for $j>1$, Equation (2.1.1) determines the first entry, the $j$th entry is $1$, and all other entries are zero. Therefore $\tau_{L}$ acts on $\ell$ by 1-1. For a general $\vec{L}\in T$, $\tau\vec{L}$ acts on the determinant line by $(-1)^{p\vec{L}}$. Since we’re assuming that $\tau\vec{L}$ acts trivially on $H_{n}(F)$, it must act trivially on $\ell$ and $2$ must divide $p\vec{L}$. ∎
@@ 2.3. Some contact-topological motivations
Now we use these contact-topological techniques to establish the existence many positive relations.
Theorem 2.3.1.
Let $n\in\mathbb{Z}_{\geq 2}$ and let $m$ be an integer which we require to be even when $n$ is odd. Then there is a Weinstein $(F,\beta_{F})$ of $\dim=2n-2$ having positive relation $\vec{L},\vec{L}^{\prime}$ with $|\vec{L}|-|\vec{L}^{\prime}|=m$.
Clearly whenever $m\neq 0$, the positive relation $\vec{L},\vec{L}^{\prime}$ cannot be braid-like. Note that the condition on the parity of $m$ cannot be removed by Lemma 2.2.1. To prove the theorem, we start with an abstract construction.
@@
Let $(X,J)$ be an almost complex manifold of $\dim X=2n$ and let $\xi_{X}=TX$ be the associated almost contact structure on $M=\mathbb{S}^{1}\times X$. By [BEM15] – which includes a definition of almost contact structures – there is a (genuine) contact structure $\xi$ on $M$ which is homotopic to $\xi_{X}$ through almost contact structures. By [HH19], we may apply a $\mathcal{C}^{0}$ small perturbation to $\{0\}\times X\subset M$ to obtain an orientable convex hypersuface $S\subset M$. Since $TS=\mathbb{R}\oplus\xi$, the restriction of $\xi|_{W}$ gives $(S,\xi)$ the structure of a stably almost complex manifold. The fact that $\xi_{X}$ and $\xi$ are homotopic implies that the cobordism class $(S,\xi)$ agrees with that of $(X,J)$ in the (stable) complex cobordism ring $\Omega^{U}_{2n}$. In particular
$\int_{S}e(\xi)=\int_{S}c_{n}(\xi)=\int_{X}c_{n}(TX)=\int_{X}e(TX)=\chi(X).$
Lemma 2.3.3.
Every class $\eta\in\Omega^{U}_{2n}$ can be represented by a connected convex hypersurface when $n\geq 2$.
Proof.
To start we note that every $\eta$ can be represented as a disjoint union $(X,J)$ of algebraic varieties, cf. [St68, §VII]. Apply the construction of the previous paragraph to such a possibly disconnected $X$, obtaining a $(S_{0},\xi_{0})$ in some (possibly disconnected) $(M_{0},\xi_{0})$ representing the bordism class $\eta$. After applying contact connected sums to the components of $(M_{0},\xi_{0})$, we obtain a connected $(M,\xi)$ and a connected convex hypersurface representative $W\subset(M,\xi)$ of the homology class $S_{0}]\in H_{2n}(M)$ will be as desired. ∎
Proof of Theorem 2.3.1.
Let $n,m$ be as in the statement of the theorem. Choose some $\eta\in\Omega^{U}_{2n}$ with $c_{n}(\eta)=m$. Explicitly, one could take
$\eta=\begin{cases}m[\mathbb{CP}^{n}]-\frac{m}{2}[\mathbb{CP}^{1}\times\mathbb{CP}^{n-1}]&n\ \text{odd},\\ m(n-1)[\mathbb{CP}^{n}]-\frac{mn}{2}[\mathbb{CP}^{1}\times\mathbb{CP}^{n-1}]&n\ \text{even}.\end{cases}$
@@
Applying the preceding lemma, we let $(S,\xi)$ be a connected convex hypersurface representing the class $\eta$. By [Br23, Theorem 1.8] there is some Weinstein $(F,\beta_{F})$ of $\dim=2n-2$ having a positive relation $\vec{L},\vec{L}^{\prime}$ determining $(W,\xi)$. By Equation (2.3.1), $\int_{S}e(\xi)=m$ determines the difference in the lengths of the positive factorizations. ∎
@@
As convex hypersurfaces $S$ are only $\mathcal{C}^{0}$-dense, rather than $\mathcal{C}^{1}$-dense in $\dim\geq 5$ contact manifolds (cf. [Ch24]), it is very difficult to explicitly describe such convex surface neighborhoods $(N\supset S,\xi)$, even for convex hypersurfaces in well-studied $(M,\xi)$. In searching for symplectic mapping class relations, we seek to have some interesting examples of convex hypersurfaces at least for which the $(F,\beta_{F})$ can be described somewhat explicitly (e.g. as affine or quasi-projective varieties). In a follow-up article, we will study the $(N,\xi)$ determined by some of the positive relations in this article using the techniques of [Av23a] and establish that each $\eta\in\Omega^{U}_{2n}$ can in fact be represented by a convex hypersurface with a tight neighborhood.
@@ 2.4. Review of algebraic Lefschetz pencils
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
The pencil $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ determines a compactly supported symplectomorphism of $(F,d\beta)$ in two ways. One way to express this symplectomorphism comes from studying the monodromy of $\pi$. Let $\eta_{i}$ be a collection of embedded, oriented, $[0,1]$-parameterized paths in $\mathbb{C}$ indexed by the $z_{i}\in\operatorname{Critv}(\pi)$, beginning at $0\in\mathbb{C}$, ending at $z_{i}$, and disjoint away from $0\in\mathbb{C}$. We require that the indices $i$ are ordered so that tangent vectors of the $\eta_{i}$ at 0 are ordered counterclockwise. Each $\eta_{i}$ determines a Lagrangian disk $D_{i}$ in $W$ for which $\pi D_{i}=\eta$ and for which $L_{i}=\partial D_{i}$ is an exact Lagrangian sphere in $F$. The $\eta_{i}$ are vanishing paths and the $D_{i}$ are their associated Lefschetz thimbles which provide framings of the $L_{i}$. Let $\gamma\subset\mathbb{C}\setminus\operatorname{Critv}(\pi)$ be a parameterized simple closed curve based at 0 and oriented counterclockwise, otherwise disjoint from the $\eta_{i}$, encircling all of the $\eta_{i}$, and encircling $\eta_{1}$ first. Then the monodromy morphism $\Phi_{\gamma}\in\operatorname{Symp}^{c}(F,\omega_{\operatorname{\mathcal{L}}})$ can be expressed as a product of Dehn twists